# Accounting for the geometry of the lung in respiratory viral infections

### v1

@@ Abstract
Increasingly, mathematical models of viral infections have come to recognise the important role of spatial structure in infection dynamics. Almost invariably, spatial models of viral infections make use of a wide, flat computational domain which is assumed to be representative of the entire affected tissue. Implicit in this assumption is that either the tissue being modelled is largely wide and homogeneous, or that the topology of the tissue has little influence on the dynamics of the system. This assumption fails to take into account the distinctive geometry of the lung. The lung is characterised by a tubular, highly branching structure, and moreover is spatially heterogeneous: deeper regions of the lung are composed of far narrower airways and are associated with more severe infection. Here, we extend a typical multicellular model of viral dynamics to account for two essential features of the geometry of the lung: the tubular structure of airways, and the branching process between airway generations. We show that, with this more realistic tissue geometry, the dynamics of infection are substantially changed compared to the standard approach, and that the resulting model is equipped to tackle important biological phenomena that are not well–addressed with existing models, including viral lineage dynamics in the lung, and heterogeneity in immune responses to infection in different regions of the respiratory tree.
Keywords: Lung, Mathematical model, Multicellular model, Viral dynamics, Influenza, SARS–CoV–2.
@@ 1 Introduction
Respiratory viruses represent a widespread and ongoing threat to global public health, both through seasonal epidemics of viruses like influenza or respiratory syncytial virus (RSV), and pandemic viruses like SARS–CoV–2 and newly emergent influenza strains. Within the host, viral infections of the lung can have highly varied patterns of progression, and can lead to marked differences in pathogenicity. In general, milder illnesses tend to arise from infections in the upper airways, such as the nasopharynx and trachea, whereas more severe and potentially life–threatening illnesses typically involve infection of the lower airways and alveolar spaces of the deep lung [20, 31, 15, 28]. Mathematical models of viral infections within the host have offered important insights into the dynamics of respiratory viruses, including how these infections spread in space [27, 15, 2]. However, traditional approaches to modelling the spatial dynamics of viral infections — including infections of the lung — do not consider the specific geometry of the tissue. Instead, models typically assume a square tissue with either periodic or no-flux boundary conditions, on the assumption that the model domain represents a small, typical patch of the overall tissue [26, 21, 29]. Implicit in this setup is the assumption that the overall tissue is reasonably wide and homogeneous, at least compared to the size of the patch. However, this is not necessarily true of the lung. The lung is an extremely structurally complex organ, comprising 24 generations of branching airways ranging in circumference from centimetres in diameter to fractions of millimetres in the alveolar ducts [9, 22]. The epithelial lining of the respiratory tree — targeted by respiratory viruses — therefore occupies an extremely large, intricately–structured surface area within a highly compact volume [9, 28]. As such, while a flat, square model tissue may offer a reasonable approximation of the structure of the upper airways, it fails to represent the extremely narrow, highly branched structure of the lower airways, which are primarily targeted by more highly pathogenic viruses [15, 20].
@@
Two recent studies have attempted to address this limitation by using mathematical models of infection spread in space which explicitly represent structural features of the lung. Chen and colleagues studied viral infection dynamics on individual airways from different generations of the lung [9]. Their model explicitly included the mucus layer lining the susceptible tissue as the principal medium in which virions spread, and incorporated experimental measurements of its velocity at different depths in the lung. Their modelling work suggested that, based on the properties of the mucus, there was a pronounced heterogeneity in the likelihood of an infection being established in different generations of the lung. Another study by Moses and coworkers compared the dynamics of SARS–CoV–2 infection on a large two–dimensional sheet of cell to the dynamics on a full, three-dimensional computational model of the entire human lung [24]. The authors found that infection on the lung geometry was notably accelerated compared to infection of the two–dimensional tissue, and resulted in an enhanced virion production. These two works hint at the heterogeneity between the dynamics of different regions of the lung, and suggest that there are important distinctions between the dynamics of infection on tissue with realistic, lung–like geometry, and more standard, flat model tissues. However, due to the complexity of both models, it is not clear how the tissue geometry influences the dynamics of the infection, or what predictive capacity this may confer to the model. There is therefore a need for a more detailed analysis of the role of the structure of the lung in spatial models of respiratory viral infections.
In this work, we seek to explore how the two main features of lung geometry — the tubular and branching tissue structure — influence the dynamics of a multicellular model of viral infection. We simulate our model on computational tissues in varying lung–like geometries to explore the role that different structural properties play in influencing infection dynamics. Our results pinpoint the dynamics possible on tissue geometries which more accurately reflect the lung, and demonstrate applications of this new model to important biological questions on the dynamics of respiratory viral infections in vivo.
@@ 2 Methods
In this work, we extend a relatively simple and established agent-based model based on one which we and others have used in earlier works [30, 5, 11]. This model couples a grid of discrete, spatially–explicit cells to a viral density surface. We assume that, at a given time, cells are in one of the following states: susceptible to infection (target), infected but not yet infectious (eclipse), productively infectious (infected), or dead (dead). Infection can arise from either cell–free virions at rate $\beta$ — secreted by productively infectious cells at rate $p$ — or from direct cell–to–cell infection between susceptible and productively infectious cells at rate $\alpha$. Latently infected cells undergo an eclipse phase of gamma–distributed duration with mean $1/\gamma$ and shape parameter $K$. In line with experimental observations for SARS–CoV–2, we assume that the cell-to–cell infection mechanism is the dominant mode of infection and tune our model parameters such that approximately 90% of infections arise from this mechanism; the other parameters are taken from fitting an equivalent ODE model to experimental data. For full details on parameter selection, refer to Supplementary Section S2. Extracellular viral density diffuses across the tissue according to linear diffusion with coefficient $D$ and uniformly decays in the environment at rate $c$. We show a schematic of this model setup in Figure 1(a). For full details of the model and its implementation, refer to the Supplementary Information.
@@
Figure 1: (a) Schematic illustrating the basic components and dynamics of the model used in this work. (b), (c) Sketch of the tube and branching geometry, respectively, as used in this work.
@@
We simulate our model in either tubular or branching geometries. In each case, the unrolled tissue is a 2D, hexagonally–packed grid of cells (consistent with the packing of real epithelial monolayers [18, 3, 5]), where the geometry is implemented in the boundary conditions of the sheet. For the tube geometry, we impose periodic boundary conditions on the top and bottom edges with zero–flux boundary conditions on the left and right edges of the tissue, which we show in Figure 1(b). Tissues with branching geometry are essentially an ensemble of tubes of varying dimensions. We assume that the tissue undergoes even, binary division at a sequence of branching points along the horizontal. Each offspring tube has periodic boundary conditions on its top and bottom edge and is assumed to be half the width (circumference) of its parent, which we show in Figure 1(c). This ensures that the overall shape of the unrolled branching tree maintains constant width. We impose zero–flux boundaries on the overall branching tree. For full details of the implementation of the tissue geometries, refer to Supplementary Sections S3 and S5.
@@ 3 Results
3.1 Diffusion of extracellular virions strongly influences viral dynamics on tubular tissues
@@
Figure 2: (a) Schematic illustrating our implementation of a tube-shaped tissue, as the surface of a cylinder of cells of circumference $C$ and length $L$. (b) The unrolled tissue depicted in (a), indicating the periodic horizontal boundaries and no-flux vertical boundaries, along with our setup for this computational experiment. We initiate infection with four infected cells (green) placed randomly on the left edge of a tube of cells (blue) with a specified aspect ratio and viral diffusion coefficient $D$. We conduct simulations for varying choices of aspect ratio and $D$, where the number of cells in the overall tissue is held fixed. (c) Snapshots of the time evolution of a typical infection in this experimental setup with aspect ratio 0.25 and $D=100~{}\text{CD}^{2}\text{h}^{-1}$. We show dead cells in black. (d) Time series of the cumulative infected cell proportion for varying $D$ (here aspect ratio is fixed at 0.25). We show ten trajectories in each case. $t_{0.95}^{D}$ denotes the average time at which simulations with viral diffusion coefficient $D$ infect 95% of the tissue. (e) Sensitivity to diffusion. For a given tissue aspect ratio, we define diffusion sensitivity as the ratio between time taken to infect 95% of the tissue with zero diffusion, $t_{0.95}^{0}$, (or $D=100~{}\text{CD}^{2}\text{h}^{-1}$, $t_{0.95}^{100}$) compared to the 95% infection time with infinite diffusion, $t_{0.95}^{\infty}$. We show diffusion sensitivity as a function of tissue aspect ratio.
@@
In Figure 2(d), we plot the cumulative proportion of infected cells in the tissue as a time series for a range of choices of the extracellular virus diffusion coefficient $D$, ranging from zero to infinity. Tissue aspect ratio is fixed here at 0.25, and we plot ten trajectories for each value of $D$. Figure 2(d) shows cumulative infected proportion growing approximately linearly with time, but with a dramatic dependence on the value of $D$. Simulations with zero diffusion (i.e., where infection can only progress through cell–to–cell infection, shown in dark green) and those where diffusion is infinite (i.e., where viral concentration is uniform, yellow) exhibit extremely different time scales of infection. Intermediate values of diffusion give rise to relatively evenly spaced trajectories between these extremes.
@@
As a metric of sensitivity to diffusion for a given choice of tissue aspect ratio, we compute the ratio between the time taken to (almost) completely infect the tissue with no diffusion and the time taken with infinite diffusion. We use the notation $t_{0.95}^{D}$ for the mean time taken to infect 95% of the cell sheet when the diffusion coefficient is $D$. We report the time until 95% infection instead of total infection due to the high degree of noise associated with the latter. In Figure 2(e) we plot $t_{0.95}^{0}/t_{0.95}^{\infty}$ for varying tissue aspect ratios in orange. Figure 2(e) shows that sensitivity to the diffusion coefficient is dramatically increased when the tissue aspect ratio shrinks, that is, when the tube becomes increasingly narrow and elongated. Figure 2(e) demonstrates that, compared to standard tissue geometries (where the aspect ratio is approximately 1), the value of the diffusion coefficient ($D$) imposes a far greater difference on system dynamics when the tissue geometry is tubular, representing the respiratory airways. We found that if the circumference and the length of the tube were equal (1:1), the infection time was approximately five–fold faster in the infinite diffusion case compared to the zero diffusion case, whereas when the aspect ratio was much smaller (1:256), this infinite diffusion case was around 80 times faster. We also compared the ratio of the time to 95% infection for the $D=100~{}\text{CD}^{2}\text{h}^{-1}$ case — which we use in the remainder of this work — compared to the infinite diffusion case, and show these results on the same axes in purple. While this curve shows closer agreement with the infinite diffusion infection time compared to the zero–diffusion case, there is nonetheless a more than six–fold difference in infection time in the most elongated geometry, and even when the aspect ratio is 0.25, the $D=100~{}\text{CD}^{2}\text{h}^{-1}$ case is still around 25% slower than the infinite diffusion case. Hence, even with this fairly large diffusion coefficient, we still obtain significantly different dynamics compared to when diffusion is infinite. As such, the assumption that viral diffusion can be ignored and approximated as infinite — sometimes justified on square tissue geometries [5, 30, 17] — does not apply to infections on more realistic, tubular geometries.
3.2 Lung branching structure imposes directionality of infection
@@
To investigate the role that the branching structure of the lung imposes on viral dynamics, we constructed a simple tissue geometry which serves as an approximation of a segment of the airway tree, and perform simulations which are presented in Figure 3. An illustration of this construction is shown in Figure 3(b). Figure 3(b) shows, from left to right, a single tube of tissue which splits into two offspring tubes at each of a sequence of branching points. The regions of tissue between branching points are the generations of the branching tree. We will make the simplifying assumption throughout this work that each tube division is even and binary — that is, if the circumference of tubes in Generation $g$ is $C_{g}$, then the circumference of tubes in Generation $g+1$ is $C_{g}/2$ — and moreover, that each generation is of equal length. We briefly note that, in reality, measurements of the anatomy of the human respiratory tree suggest that airways in subsequent generations tend to only shrink by around 20%, and as such, the sum of circumferences in subsequent generations increases substantially [22, 9]. In choosing to keep this quantity fixed, our model geometry should therefore be considered a caricature of the structure of the lung, instead of an exact representation of the biological reality. This approach, however, enables us to focus our results specifically on the presence of branching structure within the model tissue, and the presence of wider and narrower regions of tissue — both of which certainly are found in the lung — without the conflating effects of a domain which grows in overall width from left to right. Moreover, this geometry allows for the airway tree to be represented as a two-dimensional grid of cells with judiciously placed periodic boundaries, ensuring computational tractability and avoiding mathematical complications near generation splits. We show a schematic of the unrolled branching tissue in Figure 3(a), which shows the pairs of periodic boundaries lining either side of each branch of tissue within a given generation, defining the structure of the branching tree. Throughout this work, we model five tissue generations of length 100 cells each, where the circumference of the Generation 1 tube is 64 cells and the Generation 5 tubes therefore have circumference 4 cells. As with the tube, we place no–flux boundary conditions on the left and right edges of the branching tree. For full details of implementation, refer to Section 2 and the Supplementary Information.

### v4

# Accounting for the geometry of the respiratory tract in viral infections

@@ Abstract
Increasingly, experimentalists and modellers alike have come to recognise the important role of spatial structure in infection dynamics. Almost invariably, spatial computational models of viral infections — as with in vitro experimental systems — represent the tissue as wide and flat, which is often assumed to be representative of entire affected tissue within the host. However, this assumption fails to take into account the distinctive geometry of the respiratory tract in the context of viral infections. The respiratory tract is characterised by a tubular, branching structure, and moreover is spatially heterogeneous: deeper regions of the lung are composed of far narrower airways and are associated with more severe infection. Here, we extend a typical multicellular model of viral dynamics to account for two essential features of the geometry of the respiratory tract: the tubular structure of airways, and the branching process between airway generations. We show that, with this more realistic tissue geometry, the dynamics of infection are substantially changed compared to standard computational and experimental approaches, and that the resulting model is equipped to tackle important biological phenomena that do not arise in a flat host tissue, including viral lineage dynamics, and heterogeneity in immune responses to infection in different regions of the respiratory tree. Our findings suggest aspects of viral dynamics which current in vitro systems may be insufficient to describe, and points to several features of respiratory infections which can be experimentally assessed.
Keywords: Respiratory tract, Mathematical model, Multicellular model, Viral dynamics, Influenza, SARS–CoV–2.
@@ 1 Introduction
Respiratory viruses represent a widespread and ongoing threat to global public health, both through seasonal epidemics of viruses like influenza or respiratory syncytial virus (RSV), and pandemic viruses like SARS–CoV–2 and newly emergent influenza strains. Within the host, viral infections of the respiratory tract can have highly varied patterns of progression, and can lead to marked differences in pathogenicity. In general, milder illnesses tend to arise from infections in the upper airways, such as the nasopharynx and trachea, whereas more severe and potentially life–threatening illnesses typically involve infection of the lower airways and alveolar spaces of the deep lung [20, 33, 15, 30]. Mathematical models of viral infections within the host have offered important insights into the dynamics of respiratory viruses, including how these infections spread in space [29, 15, 2]. However, traditional approaches to modelling the spatial dynamics of viral infections — including infections of the respiratory tract — do not consider the specific geometry of the tissue. Instead, models typically assume a flat, roughly square tissue with either periodic or no-flux boundary conditions [27, 22, 31]. This approach mirrors that of in vitro infection assays. In both cases, these model systems (computational and experimental) often rely on the assumption that the chosen model domain represents a small, typical patch of the overall tissue within the host. Implicit in this setup is the assumption that the overall tissue is reasonably wide and homogeneous, at least compared to the size of the patch. However, this is not necessarily true of the respiratory tract. The lung is an extremely structurally complex organ, comprising 24 generations of branching airways ranging in circumference from centimetres in diameter to fractions of millimetres in the alveolar ducts [9, 23]. The epithelial lining of the overall respiratory tree — targeted by respiratory viruses — therefore occupies an extremely large, intricately–structured surface area within a highly compact volume [9, 30]. As such, while a flat, square model tissue may offer a reasonable approximation of the structure of the upper airways, it fails to represent the extremely narrow, highly branched structure of the lower airways, which are primarily targeted by more highly pathogenic viruses [15, 20].
@@
Two recent computational modelling studies have attempted to address this limitation by using models of infection spread in space which explicitly represent structural features of the respiratory tract. Chen and colleagues studied viral infection dynamics on individual airways from different generations of the lung [9]. Their model explicitly included the mucus layer lining the susceptible tissue as the principal medium in which virions spread, and incorporated experimental measurements of its velocity at different depths in the lung. Their modelling work suggested that, based on the properties of the mucus, there was a pronounced heterogeneity in the likelihood of an infection being established in different generations of the lung. Another study by Moses and coworkers compared the dynamics of SARS–CoV–2 infection on a large two–dimensional sheet of cell to the dynamics on a full, three-dimensional computational model of the entire human respiratory tract [25]. The authors found that infection on the realistic, three–dimensional geometry was notably accelerated compared to infection of the two–dimensional tissue, and resulted in an enhanced virion production. These two works hint at the heterogeneity between the dynamics of different regions of the respiratory tract, and suggest that there are important distinctions between the dynamics of infection on tissues with realistic geometry, and more standard, flat model tissues. However, due to the complexity of the models used by both sets of authors, it is not clear how the tissue geometry, specifically, influences the dynamics of the infection, or what predictive capacity may be conferred to models which account for this in a more realistic way. There is therefore a need for a more focused analysis, using models of reduced complexity, of the role of the structure of the respiratory tract in spatial models of respiratory viral infections.
In this work, we seek to explore through simulation study how the two main features of respiratory tract geometry — the tubular and branching tissue structure — influence the dynamics of a multicellular model of viral infection. We simulate our model on computational tissues with varying geometries representative of the respiratory tract to explore the role that different structural properties play in influencing infection dynamics. Our results show that the inclusion of realistic features of tissue geometry is sufficient to generate infection dynamics which are substantially different than those on flat, unstructured tissues. This suggests that the structure of the respiratory tract may have a substantial impact on the way respiratory viral infections progress within the host, and may have broader relevance to infections by other types of pathogens such as bacteria. Our findings demonstrate a clear need for improved experimental systems capable of representing the dynamics of respiratory infections in situ. Aided by the future generation of appropriate data, we show that models like ours — combined with the incorporation of more detailed descriptions of pathogen–host interactions — will be necessary to address important biological questions on the dynamics of respiratory infections in vivo.
@@ 2 Methods
In this work, we extend a relatively simple and established agent-based model based on one which we and others have used in earlier works [32, 5, 11], and which is fully described in Supplementary Section S1. This model couples a grid of discrete, spatially–explicit epithelial cells to a viral density surface. We assume that, at a given time, cells are in one of the following states: susceptible to infection (target), infected but not yet infectious (eclipse), productively infectious (infected), or dead (dead). Infection can arise from either cell–free virions at rate $\beta$ — secreted by productively infectious cells at rate $p$ — or from direct cell–to–cell infection between susceptible and productively infectious cells at rate $\alpha$. Latently infected cells undergo an eclipse phase of gamma–distributed duration with mean $1/\gamma$ and shape parameter $K$. In line with experimental observations for SARS–CoV–2, we assume that the cell-to–cell infection mechanism is the dominant mode of infection and tune our model parameters such that approximately 90% of infections arise from this mechanism; the other parameters are taken from fitting an analogous ODE model to experimental data for influenza infection in vitro. For full details on parameter selection, refer to Supplementary Section S2. Extracellular viral density diffuses across the tissue according to linear diffusion with coefficient $D$ and uniformly decays in the environment at rate $c$. This model and its parameterisation were developed with SARS–CoV–2 and influenza in mind, but are readily generalisable to other respiratory viruses. We show a schematic of this model setup in Figure 1(a). For full details of the model and its implementation, refer to the Supplementary Information.
@@
Figure 1: (a) Schematic illustrating the basic components and dynamics of the model used in this work. (b), (c) Sketch of the tube and branching geometry, respectively, as used in this work. Pairs of dashed lines of the same colour indicate pairs of periodic boundaries, while solid black lines in (c) represents cuts in the sheet.
@@
We simulate our model in either tubular or branching geometries. In each case, the unrolled tissue is a 2D, hexagonally–packed grid of cells (consistent with the packing of real epithelial monolayers [18, 3, 5]), where the geometry is implemented in the boundary conditions of the sheet. For the tube geometry, we impose periodic boundary conditions on the top and bottom edges with no-flux boundary conditions on the left and right edges of the tissue, which we show in Figure 1(b). No-flux (as opposed to periodic) boundary conditions are imposed on the ends of the tube to ensure that — in keeping with the biological reality — infection cannot jump from one end of the tube to the other. This choice of boundary conditions, moreover, prevents any viral density from leaving the system.
Tissues with branching geometry are essentially an ensemble of tubes of varying dimensions. We assume that the tissue undergoes even, binary division at a sequence of branching points along the horizontal. Each offspring tube has periodic boundary conditions on its top and bottom edge and is assumed to be half the width (circumference) of its parent, which we show in Figure 1(c). This ensures that the overall shape of the unrolled branching tree maintains constant width. We impose zero–flux boundaries on the ends of the overall branching tree. For full details of the implementation of the tissue geometries, refer to Supplementary Sections S3–S5.
@@ 3 Results
3.1 Dependence of viral dynamics on the rate of viral diffusion is increased on tubular tissues
@@
Figure 2: (a) Schematic illustrating our implementation of a tube-shaped tissue, as the surface of a cylinder of cells of circumference $C$ and length $L$. (b) The unrolled tissue depicted in (a), indicating the periodic horizontal boundaries and no-flux vertical boundaries, along with our setup for this computational experiment. We initiate infection with four infected cells (green) placed randomly on the left edge of a tube of cells (blue) with a specified aspect ratio ($C/L$) and viral diffusion coefficient $D$. We conduct simulations for varying choices of aspect ratio and $D$, where the number of cells in the overall tissue is held fixed. (c) Snapshots of the time evolution of a typical infection in this experimental setup with aspect ratio 0.25 and $D=100~{}\text{CD}^{2}\text{h}^{-1}$. We show dead cells in black. (d) Time series of the cumulative infected cell proportion for varying $D$ (here aspect ratio is fixed at 0.25). We show ten trajectories in each case. $t_{0.95}^{D}$ denotes the average time at which simulations with viral diffusion coefficient $D$ infect 95% of the tissue. (e) Sensitivity to diffusion. For a given tissue aspect ratio, we define diffusion sensitivity as the ratio between time taken to infect 95% of the tissue with zero diffusion, $t_{0.95}^{0}$, (or $D=100~{}\text{CD}^{2}\text{h}^{-1}$, $t_{0.95}^{100}$) compared to the 95% infection time with infinite diffusion, $t_{0.95}^{\infty}$. We show diffusion sensitivity as a function of tissue aspect ratio.
@@
In Figure 2(d), we plot the cumulative proportion of infected cells in the tissue as a time series for a range of choices of the extracellular virus diffusion coefficient $D$, ranging from zero to infinity. Tissue aspect ratio is fixed here at 0.25, and we plot ten trajectories for each value of $D$. Figure 2(d) shows cumulative infected proportion growing approximately linearly with time, but with a dramatic dependence on the value of $D$. Simulations with zero diffusion (i.e., where infection can only progress through cell–to–cell infection, shown in dark green) and those where diffusion is infinite (i.e., where viral concentration is uniform, yellow) exhibit very different time scales of infection. Intermediate values of diffusion give rise to relatively evenly spaced trajectories between these extremes.
@@
To investigate sensitivity to diffusion for a given choice of tissue aspect ratio, we compute the ratio between the time taken to (almost) completely infect the tissue with no diffusion and the time taken with infinite diffusion. We use the notation $t_{0.95}^{D}$ for the mean time taken to infect 95% of the cell sheet when the diffusion coefficient is $D$. We report the time until 95% infection instead of total infection due to the high degree of noise associated with the latter. In Figure 2(e) we plot $t_{0.95}^{0}/t_{0.95}^{\infty}$ for varying tissue aspect ratios in orange. Figure 2(e) shows that infinite viral diffusion results in a faster 95% infection time than in simulations with finite or zero viral diffusion, but that the extent to which these infection times differ is highly dependent on the tissue aspect ratio. We found that if the circumference and the length of the tube were equal (an aspect ratio of 1), the infection time was increased around five–fold in the zero diffusion case compared to the infinite diffusion case, whereas when the aspect ratio was much smaller (1/256), the zero diffusion case took 80 times longer. This is because, on narrow tubes, infections which can only spread cell–to–cell (that is, with zero viral diffusion) are very slow due to the limited number of susceptible cells neighbouring the infection front at a given time. The system is therefore particularly sensitive to the presence of viral diffusion, since cell–free infections away from the infection front — even if rare — dramatically increase access to neighbouring susceptible cells. This effect is much less pronounced on wide tubes, where there is less limitation on susceptible cells neighbouring the infection front.
In Figure 2(e), we also compare the ratio of the time to 95% infection for the $D=100~{}\text{CD}^{2}\text{h}^{-1}$ case — which we use in the remainder of this work — relative to the infinite diffusion case, and show these results on the same axes in purple. While this curve shows infection times for the infinite diffusion case are in closer agreement with the $D=100~{}\text{CD}^{2}\text{h}^{-1}$ case than the zero diffusion case, there is nonetheless a more than six–fold difference in infection time in the most elongated geometry, and even when the aspect ratio is 0.25, the $D=100~{}\text{CD}^{2}\text{h}^{-1}$ case is still around 25% slower than the infinite diffusion case. Hence, even with this fairly large diffusion coefficient, we still obtain significantly different dynamics compared to when diffusion is infinite. Notably, the change in dynamics observed in Figure 2(e) is a direct consequence of the interplay between viral diffusion and tissue geometry, and not due to a change in the prevalence of the cell–to–cell versus cell–free infection. In Supplementary Figure S2, we plot the mean proportion of infections arising from the cell–to–cell mechanism across ten simulations for a range of viral diffusion coefficients and tissue aspect ratios. Supplementary Figure S2 shows that there is no discernable difference in the balance of the two modes of transmission for any $D\geq 10~{}\text{CD}^{2}\text{h}^{-1}$ in any given tissue aspect ratio, and moreover that there was negligible variation in this balance between tissue aspect ratios relative to the noise of the system. Collectively, Figure 2(e) shows that, on increasingly narrow tube-shaped tissues (that is, tubes with small aspect ratio), the dynamics of viral infections become increasingly dependent on not just the presence, but also the rate of viral diffusion. These results suggest that the assumption that viral diffusion can be ignored and approximated as infinite — as is sometimes applied in models which use square tissue geometries [5, 32, 17] — may require greater scrutiny in tissues where infection primarily spreads along one axis and is constrained along another, as is the case for respiratory infections in vivo.
3.2 Tissue branching structure imposes directionality of infection
@@
To investigate the role that the branching structure of the respiratory tract imposes on viral dynamics, we constructed a simple tissue geometry which serves as an approximation of a segment of the airway tree, and perform simulations which are presented in Figure 3. An illustration of this construction is shown in Figure 3(b). Figure 3(b) shows, from left to right, a single tube of tissue which splits into two offspring tubes at each of a sequence of branching points. The regions of tissue between branching points are the generations of the branching tree. We will make the simplifying assumption throughout this work that each tube division is even and binary — that is, if the circumference of tubes in Generation $g$ is $C_{g}$, then the circumference of tubes in Generation $g+1$ is $C_{g}/2$ — and moreover, that each generation is of equal length. We briefly note that, in reality, measurements of the anatomy of the human respiratory tree suggest that airways in subsequent generations tend to only shrink by around 20%, and as such, the sum of circumferences in subsequent generations increases substantially [23, 9]. In choosing to keep this quantity fixed, our model geometry should therefore be considered a caricature of the structure of the respiratory tract, instead of an exact representation of the biological reality. This approach, however, enables us to focus our results specifically on the presence of branching structure within the model tissue, and the presence of wider and narrower regions of tissue — both of which certainly are found in the respiratory tract — without the conflating effects of a domain which grows in overall width from left to right. Moreover, this geometry allows for the airway tree to be represented as a two-dimensional grid of cells with judiciously placed periodic boundaries, ensuring computational tractability and avoiding mathematical complications near generation splits. We show a schematic of the unrolled branching tissue in Figure 3(a), which shows the pairs of periodic boundaries lining either side of each branch of tissue within a given generation, defining the structure of the branching tree. Throughout this work, we model five tissue generations of length 100 cells each, where the circumference of the Generation 1 tube is 64 cells and the Generation 5 tubes therefore have circumference 4 cells (again, in reality, branches of the human respiratory tract are typically wider in absolute circumference; we use this smaller size to maintain computational tractability while qualitatively representing the in vivo structure). As with the tube, we place no–flux boundary conditions on the left and right edges of the branching tree. For full details of implementation, refer to Section 2 and the Supplementary Information.